OpenStudy (anonymous):

The Sanfords' car traveled 384 miles using 12 gallons of gasoline. At this rate, how far could the car travel on a full tank of 18 gallons of gasoline? a 390 b 414 c 576 d 640

OpenStudy (anonymous):

the answer would be c

OpenStudy (shamim):

By 1 galon the car will go 384/12 Right?

OpenStudy (anonymous):

all you need to do is divide the miles given by the amount of gas used to travel that distance then multiply by the gallons desired

OpenStudy (texaschic101):

proportion time 384/12 = x/18 cross multiply 12x = 6912 x = 6912/12 x = 576

OpenStudy (shamim):

By 18 galon the car will go 384*18/12=?

OpenStudy (anonymous):

thank you texaschic101, that is a great rendition of what to do

OpenStudy (texaschic101):

it can also be done the way shamin said

OpenStudy (texaschic101):

oops..spelled name wrong...shamim